# High-Order Hermite Optimization: Fast and Exact Gradient Computation in Open-Loop Quantum Optimal Control using a Discrete Adjoint Approach

### v2

@@ Abstract
This work introduces the High-Order Hermite Optimization (HOHO) method, an open-loop discrete adjoint method for quantum optimal control. Our method is the first of its kind to efficiently compute exact (discrete) gradients when using continuous, parameterized control pulses while solving the forward equations (e.g. Schrödinger’s equation or the Linblad master equation) with an arbitrarily high-order Hermite Runge-Kutta method. The HOHO method is implemented in QuantumGateDesign.jl an open-source software package for the Julia programming language, which we use to perform numerical experiments comparing the method to Juqbox.jl. For realistic model problems we observe speedups up to 775x.
@@ 1 Introduction
QOC methods may be open-loop, in which case the quantum system is simulated numerically, or closed-loop, in which case feedback from real experiments is used. For the problem of gate design, closed-loop typically means applying control pulses and measuring the performance of the gate at the end of the pulse application (e.g. by randomized benchmarking [27] or quantum state tomography [20]). In general, open-loop methods excel at accurately computing gradients of the objective function with a cost that does not scale with the number of control parameters, whereas closed-loop methods excel at correctly characterizing the behavior of the the gate by using experimental results instead of simulated results based on potentially inaccurate theoretical models, and at scaling with the number of qubits. Control pulses found using open-loop control may be further tuned using closed-loop control, so even when closed-loop control is necessary for the final implementation of a gate, open-loop control is still a valuable tool.
The adjoint method is a technique for efficiently calculating the gradient of an objective functional that depends on the solution of a differential equation [13]. By “efficiently”, we mean in particular that the computational cost comes mainly from solving two initial value problems (the forward and adjoint equations), and does not depend significantly on the number of control parameters. Within this framework, there are continuous adjoint methods, in which the adjoint equations are derived from the continuous forward forward equations (optimize then discretize), and discrete adjoint methods, in which the adjoint equations are derived from the discretization of the forward equations by a numerical method (discretize then optimize), so that the gradients computed are exact with respect to the numerical method used. This is especially important when optimizing the objective functional using quasi-Newton methods such as L-BFGS [30], because even modest errors in the gradient may ruin the Hessian approximation, so that the advantage of using curvature information is lost.
@@
The most ubiquitous methods for open-loop QOC are Gradient Ascent Pulse Engineering (GRAPE) [11] and Krotov’s method [12]. These methods rely on the assumption that the control pulses are piecewise-constant, which allows the forward equations for be solved using matrix exponentiation for each time interval over which the control pulses are constant. Consequently, the discretized forward equations are analytically simple, which makes it easy to derive exact expressions for partial derivatives of the objective function. GRAPE is a discrete adjoint method, and updates all the control parameters at once using the gradient, while Krotov’s method updates only a single control parameter at a time, with the advantage that the update scheme is guaranteed to converge monotonically.
@@
Before GRAFS, The idea of parameterizing the control pulses in terms of basis functions was also explored in the Chopped Random Basis (CRAB) method [6], which uses smooth, parameterized control pulses. But as a tradeoff lacks a way to compute the gradient (except by finite-difference approximation), and instead uses a direct search to optimize the objective function. The Gradient Optimization of Analytic Controls (GOAT) method [26] also uses smooth, parameterized control pulses, but performs a gradient-based search by differentiating the forward equation with respect to each control parameter (forward-mode differentiation). The GOAT method produces exact gradients, but must solve an initial value problem for every control parameter, so that the gradient computation is not efficient. We also mention the Padé Integrator COllocation method (PICO) [38] which optimizes over the quantum states directly (a direct method), implementing the Schrödinger’s equation as a constraint in the optimization instead of having it enforced implicitly by a numerical timestepping scheme.
@@
In particular, reverse-mode AD must store the expression graph and the values of intermediate variables in that graph. This results in memory storage that scales the system size and with the number of matrix-vector multiplications (which scales with the number of timesteps taken, and hence also with the stiffness of the system) and makes optimal control of large quantum systems, especially open quantum systems, infeasible. In contrast, discrete adjoint methods are often able to store intermediate in a small number of pre-allocated arrays, drastically reducing the total memory required. This also drastically reduces the number of large (the size of the system) dynamic memory allocations required, which can have a significant impact on performance when solving sparse differential equations (which typically appear in QOC), since the costs of the memory allocation and iteratively solving a linear system of equations (e.g. in an implicit method) at each timestep may both be $\mathcal{O}(N)$, where $N$ is the size of the system. Additionally, the discrete adjoint method differentiates each discretized forward equation to produce a corresponding adjoint equation. This allows each both the forward and adjoint equations to be solved iteratively. In contrast, in most software implementations of AD will differentiate through each step of the iterative solver, resulting in increased memory usage and unnecessary computations compared to the discrete adjoint method.
Some of the shortcomings of AD are addressed by the semi-automatic differentiation approach [14]. In this approach, the objective functional is rewritten in terms of intermediary variables. By carefully choosing the intermediate variables the computational graph for can be kept small, reducing the total memory required compared to automatic differentiation, while still keeping the flexibility in the choice of the objective functional.
@@ 2.1 Quantum Computing
The state of a closed, quantum system is represented by a complex-valued state vector [^3] In quantum computing literature, states are often represented using bra ket notation, in which case the state $\boldsymbol{\psi}\in\mathbb{C}^{N}$ corresponds to $|\psi\rangle$ in a particular basis. Note that $|\psi\rangle$ only represents a physical state: it does not choose a basis to use to represent that state. We switch between the two notations when convenient, but mostly stick to representing states as complex-valued vectors. $\boldsymbol{\psi}(t)\in\mathbb{C}^{N}$. Each observable (a quantity which can be physically measured) of the system is associated with a linear operator. The eigenvalues of the operator are the possible outcomes of the measurement, and the corresponding eigenvectors are the states for which that measurement will be observed.
@@ 2.2 Governing Equation: Schrödinger’s Equation
A collection of $Q$ (possibly entangled) qubits is described by a state vector $\boldsymbol{\psi}(t)\in\mathbb{C}^{2^{Q}}$. The basis elements may be constructed using Kronecker products of the state vectors of each qubit. E.g. the state corresponding to a 2 qubit system with the first qubit in the ground state and the second qubit in the excited state is $|01\rangle={\mathbf{e}}_{1}\otimes{\mathbf{e}}_{2}\in\mathbb{C}^{2^{2}}$. The standard ordering of the basis elements in bra-ket notation is $|00\rangle,|01\rangle,|10\rangle,|11\rangle$. A superposition of observable states is represented by a linear combination of the standard basis elements. For example $(|0\rangle+|1\rangle)/\sqrt{2}$ corresponds to a qubit in a superposition of the ground and excited states. Upon measurement, there is a 50/50 chance of observing the system to be in either state.
@@
, and the raising operator is given by $a^{\dagger}$. For a system of $Q$ qudits, with the $q$-th qudit having $n_{q}$ levels, the lowering operator of the subsystem corresponding to the $q$-th qudit is given by
$a_{q}\coloneqq I_{n_{Q}}\otimes\cdots\otimes I_{n_{q}+1}\otimes a_{q}\otimes I_{n_{q}-1}\otimes\cdots\otimes I_{n_{1}},\quad N=\prod_{q=1}^{Q}n_{q}.$
As a representative example, for a system of $Q$ transmon qubits in the dispersive limit the drift Hamiltonian is
$H_{d}=\sum_{q=1}^{Q}\left(\omega_{q}a_{q}^{\dagger}a_{q}-\frac{\xi_{q}}{2}a_{q}^{\dagger}a_{q}^{\dagger}a_{q}a_{q}-\sum_{p>q}\xi_{pq}a_{p}^{\dagger}a_{p}a_{q}^{\dagger}a_{q}\right).$ (4)
@@ 2.3 Optimal Control Problem
More precisely, we control the time evolution of $E$ state vectors [^5] For $n$ qubits, $E=2^{n}$. When each qubit is modeled as a pure two-level system, then $N=E$. If additional levels are modeled, then $N>E$. that form a basis of the computational subspace:
@@
$\mathcal{F}(U(T;\boldsymbol{\theta}),U_{\text{target}})=\mathcal{F}(e^{i\delta_{1}}U(T;\boldsymbol{\theta}),e^{i\delta_{2}}U_{\text{target}}),\quad\ \forall\ \delta_{1},\delta_{2}\in\mathbb{R}.$
@@
It is common for optimization to expect an optimization problem to be formulated as a minimization problem, so instead of maximizing the fidelity, we often consider the equivalent problem of minimizing the trace infidelity
@@
To illustrate the idea of gate design, we provide the example of implementing Hadamard gate for a system described by the Hamiltonian
@@
which equals the standard infidelity when $U(T;\boldsymbol{\theta})$ is unitary, but is bounded between 0 and 1 even for non-unitary $U(T;\boldsymbol{\theta})$. Increasing the spectral norm of $U(T;\boldsymbol{\theta})$ increases $\left\lVert U(T;\boldsymbol{\theta})\right\rVert_{F}^{2}$ (because $\left\lVert A\right\rVert_{2}\leq\left\lVert A\right\rVert_{F}$), which deters an optimizer from exploiting this non-physical phenomenon to minimize the infidelity.
We use either the trace infidelity or the generalized infidelity as the main component of our objective function, but we may add additional terms in order to achieve other desirable properties. In our case, we also add a term measuring the population of the “guard” states (which are not in the computational subspace) throughout the duration of the gate:
[…]
Here $W$ is an $N\times N$ matrix whose null space is the computational subspace. This is done to motivate the truncation of the Hilbert space, which for most quantum computing architectures has infinite dimensions. By including the guard penalty (9) in our objective function, we prevent “leakage” of the state vector into the non-computational subspace. Physically, this means punishing the population of highly excited states. In our models, we typically include one or two guard levels for each qubit. Then, in the computational basis, $W$ is diagonal, with $W_{i,i}=1$ when the $i$-th basis state is an essential state. For a non-essential state $j$, we have $W_{j,j}=w_{j}\neq 0$, where the weight $w_{j}$ can be adjusted to suppress the population of certain guard levels more than others.
@@ 2.4 Effect of State Errors on Fidelity
The error, $\mathcal{E}(\delta,\Delta t)$ in the state vector, caused by a perturbation $\delta$ in a coefficient of the control, or the truncation error $\Delta t^{p}$ for a $p$-th order accurate numerical method, will be $\mathcal{E}(\delta,\Delta t)=\mathcal{O}(C\delta+\tilde{C}\Delta t^{p})$. Here $C$ and $\tilde{C}$ are constants that can depend on the solution and the choice of numerical method but not on $\delta$ and $\Delta t$. Here the goal is to determine the control accurately and it is therefore important that $\tilde{C}\Delta t^{p}\ll C\delta+\tilde{C}$, so that the optimization landscape depends more heavily on the real, physical change in the state vector due to the perturbation $\delta$ than on the non-physical change due to the truncation error. Clearly this is easier to achieve for a high order method.
We note, however, that the error in the infidelity is quadratic in the state error. As a result, the state error can generally be allowed to be quadratically larger than the error in the fidelity. For other objectives, the error may be linear in the state error, and hence a higher level of accuracy in the state will be required.
@@
As an example, we optimize of the generalized infidelity of a CNOT gate for a system of two qubits coupled to a resonator bus (which will be described in more detail in Section 4.2). The optimization is performed using a fourth-order method to solve Schrödinger’s equation numerically, with the timestep size chosen relative final state error is approximately $10^{-1}$. For each optimization iteration, we solve the same equation again, choosing the timestep size so that the relative final state error is approximately $10^{-7}$ In Figure 2, we plot the generalized infidelity as computed by the low-accuracy and high-accuracy solutions, as well as the relative final state error in the low-accuracy solution, as a function of the number of optimization iterations. We see that although after 500 iterations the optimization produces a gate with a generalized infidelity of approximately $10^{-4}$ as computed by the low-accuracy method, the actual generalized infidelity is closer to $10^{-1}$, and the optimization iterations performed after the generalized infidelity had reached $10^{-2}$ were wasted computations which actually increased the generalized infidelity.
Figure 2: Generalized infidelity of a CNOT gate, as computed by a low-accuracy and high-accuracy numerical method as a function of the number of optimization iterations, where the optimization is performed using the low-accuracy method.

### v3

@@ Abstract
This work introduces the High-Order Hermite Optimization (HOHO) method, an open-loop discrete adjoint method for quantum optimal control. Our method is the first of its kind to efficiently compute exact (discrete) gradients when using continuous, parameterized control pulses while solving the forward equations (e.g. Schrödinger’s equation or the Linblad master equation) with an arbitrarily high-order Hermite Runge-Kutta method. The HOHO method is implemented in QuantumGateDesign.jl, an open-source software package for the Julia programming language, which we use to perform numerical experiments comparing the method to Juqbox.jl. For realistic model problems we observe speedups up to 775x.
@@ 1 Introduction
QOC methods may be open-loop, in which case the quantum system is simulated numerically, or closed-loop, in which case feedback from real experiments is used. For the problem of gate design, closed-loop typically means applying control pulses and measuring the performance of the gate at the end of the pulse application (e.g. by randomized benchmarking [35] or quantum state tomography [26]). In general, open-loop methods excel at accurately computing gradients of the objective function with a cost that does not scale with the number of control parameters, whereas closed-loop methods excel at correctly characterizing the behavior of the gate by using experimental results instead of simulated results based on potentially inaccurate theoretical models, and at scaling with the number of qubits. Control pulses found using open-loop control may be further tuned using closed-loop control, so even when closed-loop control is necessary for the final implementation of a gate, open-loop control is still a valuable tool.
The adjoint method is a technique for efficiently calculating the gradient of an objective functional that depends on the solution of a differential equation [17]. By “efficiently,” we mean in particular that the computational cost comes mainly from solving two initial value problems (the forward and adjoint equations), and does not depend significantly on the number of control parameters. Within this framework, there are continuous adjoint methods, in which the adjoint equations are derived from the continuous forward forward equations (optimize then discretize), and discrete adjoint methods, in which the adjoint equations are derived from the discretization of the forward equations by a numerical method (discretize then optimize), so that the gradients computed are exact with respect to the numerical method used. This is especially important when optimizing the objective functional using quasi-Newton methods such as L-BFGS [39] because moderate errors in the gradient can obscure the approximate Hessians used by quasi-Newton methods [13], which results in the advantage of using curvature information being lost.
@@
The most ubiquitous methods for open-loop QOC are Gradient Ascent Pulse Engineering (GRAPE) [27, 13] and Krotov’s method [15]. These methods rely on the assumption that the control pulses are piecewise-constant, which allows the forward equations for be solved using matrix exponentiation for each time interval over which the control pulses are constant. Consequently, the discretized forward equations are analytically simple, which makes it easy to derive exact expressions for partial derivatives of the objective function. GRAPE is a discrete adjoint method, and updates all the control parameters at once using the gradient, while Krotov’s method updates only a single control parameter at a time, with the advantage that the update scheme is guaranteed to converge monotonically.
@@
Before GRAFS, The idea of parameterizing the control pulses in terms of basis functions was also explored in the Chopped Random Basis (CRAB) method [7], which uses smooth, parameterized control pulses. But as a tradeoff it lacks a way to compute the gradient (except by finite-difference approximation), and instead uses a derivative-free (zeroth-order) method [8] to optimize the objective function. The Gradient Optimization of Analytic Controls (GOAT) method [34] also uses smooth, parameterized control pulses, but performs a gradient-based search by differentiating the forward equation with respect to each control parameter (forward-mode differentiation). The GOAT method produces exact gradients, but must solve an initial value problem for every control parameter, so that the gradient computation is not efficient. We also mention the Padé Integrator COllocation method (PICO) [49] which optimizes over the quantum states directly (a direct method), implementing Schrödinger’s equation as a constraint in the optimization instead of having it enforced implicitly by a numerical timestepping scheme.
@@
In particular, reverse-mode AD must store the expression graph and the values of intermediate variables in that graph. This results in memory storage that scales with the system size and with the number of matrix-vector multiplications (which scales with the number of timesteps taken, and hence also with the stiffness of the system) and makes optimal control of large quantum systems, especially open quantum systems, infeasible. In contrast, discrete adjoint methods are often able to store intermediate variables in a small number of pre-allocated arrays, drastically reducing the total memory required. This also drastically reduces the number of large (the size of the system) dynamic memory allocations required, which can have a significant impact on performance when solving sparse differential equations (which typically appear in QOC), since the costs of allocating memory and iteratively solving a linear system of equations (e.g. in an implicit method) at each timestep are both $\mathcal{O}(N)$, where $N$ is the size of the complex-valued matrix system. For a system of $N_{Q}$ qubits, we have $N=2^{N_{Q}}$.
When the solution method for the forward equations involves solving linear systems, for example when using an implicit timestep method, the discrete adjoint method supports using iterative linear solvers such as GMRES [44] more efficiently than AD does. In particular, when solving the forward equations using an iterative linear solver, naive implementations of AD will directly differentiate through the low-level implementation of the solver [36, 25], which results in increased memory usage and unnecessary computations.
Some of the shortcomings of AD are addressed by the semi-automatic differentiation approach [18]. In this approach, the objective functional is rewritten in terms of intermediate variables. By choosing the intermediate variables carefully, the computational graph for can be kept small, reducing the total memory required compared to automatic differentiation, while still keeping the flexibility in the choice of the objective functional.
@@
In principle, Störmer-Verlet and the implicit midpoint method can be made higher-order using compositional methods [24, 30], which divide each timestep into a sequence of forward and backward steps. However, the number of stages in compositional methods grows rapidly, with the tenth-order method requiring 35 stages. Thus if the base timestep method requires solving a linear system of equations (which is true of both the implicit midpoint method and Störmer-Verlet when applied to Schrödinger’s equation), then the number of linear solves required to take a full timestep using the compositional method will be much larger than for HOHO. The linear solve is typically the most time consuming part of the solver. Furthermore, since the compositional methods requiring taking steps backward in time, it is unclear if they can be stably applied to problems with damping, which can occur in open quantum systems.
@@ 2.1 Quantum Computing
The state of a closed, quantum system is represented by a complex-valued state vector [^3] In quantum computing literature, states are often represented using bra-ket notation, in which case the state $\boldsymbol{\psi}\in\mathbb{C}^{N}$ corresponds to $|\psi\rangle$ in a particular basis. Note that $|\psi\rangle$ only represents a physical state: it does not choose a basis to use to represent that state. We switch between the two notations when convenient, but mostly stick to representing states as complex-valued vectors. $\boldsymbol{\psi}(t)\in\mathbb{C}^{N}$. Each observable (a quantity which can be physically measured) of the system is associated with a linear operator. The eigenvalues of the operator are the possible outcomes of the measurement, and the corresponding eigenvectors are the states for which that measurement will be observed.
@@ 2.2 Governing Equation: Schrödinger’s Equation
A collection of $N_{Q}$ (possibly entangled) qubits is described by a state vector $\boldsymbol{\psi}(t)\in\mathbb{C}^{2^{{N_{Q}}}}$. The basis elements may be constructed using Kronecker products of the state vectors of each qubit. E.g. the state corresponding to a 2 qubit system with the first qubit in the ground state and the second qubit in the excited state is $|01\rangle={\mathbf{e}}_{1}\otimes{\mathbf{e}}_{2}\in\mathbb{C}^{2^{2}}$. The standard ordering of the basis elements in bra-ket notation is $|00\rangle,|01\rangle,|10\rangle,|11\rangle$. A superposition of observable states is represented by a linear combination of the standard basis elements. For example $(|0\rangle+|1\rangle)/\sqrt{2}$ corresponds to a qubit in a superposition of the ground and excited states. Upon measurement, there is a 50/50 chance of observing the system to be in either state.
@@
and the raising operator is given by $a^{\dagger}$. For a system of $N_{Q}$} qudits, with the $q$-th qudit having $n_{q}$ levels, the lowering operator of the subsystem corresponding to the $q$-th qudit is given by
$a_{q}\coloneqq I_{n_{{N_{Q}}}}\otimes\cdots\otimes I_{n_{q}+1}\otimes a_{q}\otimes I_{n_{q}-1}\otimes\cdots\otimes I_{n_{1}},\quad N=\prod_{q=1}^{{N_{Q}}}n_{q}.$
As a representative example, for a system of $N_{Q}$} transmon qubits in the dispersive limit the drift Hamiltonian is
$H_{d}=\sum_{q=1}^{{N_{Q}}}\left(\omega_{q}a_{q}^{\dagger}a_{q}-\frac{\xi_{q}}{2}a_{q}^{\dagger}a_{q}^{\dagger}a_{q}a_{q}-\sum_{p>q}\xi_{pq}a_{p}^{\dagger}a_{p}a_{q}^{\dagger}a_{q}\right).$ (4)
@@ 2.3 Optimal Control Problem
More precisely, we control the time evolution of $E$ state vectors [^5] For $n$ qubits, $E=2^{n}$. If each qubit is modeled as a pure two-level system, then $N=E$. If additional levels are modeled, then $N>E$. that form a basis of the possible initial states that the gate may be applied to:
@@
$\mathcal{F}(U(T;\boldsymbol{\theta}),U_{\text{target}})=\mathcal{F}(e^{i{\phi}_{1}}U(T;\boldsymbol{\theta}),e^{i{\phi}_{2}}U_{\text{target}}),\quad\ \forall\ {\phi}_{1},{\phi}_{2}\in\mathbb{R}.$
@@
It is common for optimization software to expect an optimization problem to be formulated as a minimization problem, so instead of maximizing the fidelity, we often consider the equivalent problem of minimizing the trace infidelity
@@
To illustrate the idea of gate design, we provide the example of implementing a Hadamard gate for a system described by the Hamiltonian
@@
which equals the standard infidelity when $U(T;\boldsymbol{\theta})$ is unitary, but is non-negative even for non-unitary $U(T;\boldsymbol{\theta})$. Increasing the spectral norm of $U(T;\boldsymbol{\theta})$ increases $\left\lVert U(T;\boldsymbol{\theta})\right\rVert_{F}^{2}$ (because $\left\lVert A\right\rVert_{2}\leq\left\lVert A\right\rVert_{F}$), which deters an optimizer from exploiting this non-physical phenomenon to minimize the infidelity.
We emphasize that our main motivation for using the generalized infidelity is not to force numerical approximations of $U(T;\boldsymbol{\theta})$ to be unitary even when the numerical approximation is inaccurate, but rather to avoid misleading objective function values that are near zero (or worse, negative) even though $U(T;\boldsymbol{\theta})$ and $U_{\textrm{target}}$ are not closely aligned. This is especially useful for choosing stopping criterion based on the objective function value, for example choosing to stop the optimization once the (generalized) infidelity falls below a certain threshold, indicating that the gate has reached a desired level of accuracy. In general, the extent to which a numerical approximation of $U(t;\boldsymbol{\theta})$ is unitary throughout $t\in[0,T]$ depends on the accuracy of the approximation. Later on in the numerical experiments section of this paper (Section 4.2) we study this. The result of that experiment is displayed in Table 1, which shows the convergence of the unitarity of the numerical approximation of $U(t;\boldsymbol{\theta})$ as the accuracy of a numerical method increases.
Order 2 Order 4 Order 6 Order 8 Order 10
Steps Err Cvg Err Cvg Err Cvg Err Cvg Err Cvg
$^{10} 2.2$(-2) -4.0 6.3(-3) 1.2 3.7(-5) 7.3 4.0(-7) 7.8 2.8(-9) 9.8
$2^{11}$ 1.7(-2) 3.6(-1) 1.4(-4) 5.5 6.1(-7) 5.9 1.6(-9) 7.9 2.8(-12) 9.9
$2^{12}$ 8.3(-3) 1.1 8.6(-6) 4.1 9.7(-9) 6.0 6.6(-12) 8.0 1.9(-13) 3.9
$2^{13}$ 1.4(-3) 2.6 5.4(-7) 4.0 1.5(-10) 6.0 1.7(-13) 5.3 1.8(-13) 3.4(-2)
$2^{14}$ 2.8(-4) 2.3 3.4(-8) 4.0 2.5(-12) 5.9 2.2(-13) -3.8(-1) 2.0(-13) -1.6(-1)
$2^{15}$ 6.7(-5) 2.1 2.1(-9) 4.0 2.3(-12) 1.1(-1) 2.2(-12) -3.3 2.1(-12) -3.4
Table 1: Error and convergence rates of the unitarity of the state, as measured by the discrete $L^{2}$ norm of $\frac{1}{\sqrt{E}}\|U(t;\boldsymbol{\theta})\|_{F}$ across all timesteps, for the three-subsystem example when using the “best” control for implementing a CNOT gate in the two qudits plus resonator model (see Figure 5), using the Hermite method with orders two through ten.
We use either the trace infidelity or the generalized infidelity as the main component of our objective function, but we may add additional terms in order to achieve other desirable properties. In our case, we also add a term measuring the population of the “guard” states (which are not in the essential subspace) throughout the duration of the gate:
$\mathcal{G}(U(\cdot;\boldsymbol{\theta}))=\frac{1}{T}\int_{0}^{T}\left\langle U(t;\boldsymbol{\theta}),WU(t;\boldsymbol{\theta})\right\rangle_{F}dt.$ (9)
Here $W$ is an $N\times N$ matrix whose null space is the essential subspace. This is done to motivate the truncation of the Hilbert space, which for most quantum computing architectures has an infinite number of dimensions. By including the guard penalty (9) in our objective function, we prevent “leakage” of the state vector into the non-essential (also called the guard subspace), which is not used for computation. Physically, this means punishing the population of highly excited states. In our models, we typically include one or two guard levels for each qubit. Then, in the computational basis, $W$ is diagonal, with $W_{i,i}=0$ when the $i$-th basis state is an essential state. For a non-essential state $j$, we have $W_{j,j}=w_{j}\neq 0$, where the weight $w_{j}$ can be adjusted to suppress the population of certain guard levels more than others.
@@ 2.4 Effect of State Errors on Fidelity
The error $\mathcal{E}(\delta,\Delta t)$ in the state vector caused by a perturbation $\delta$ in a coefficient of the control and the truncation error $\Delta t^{p}$ for a $p$-th order accurate numerical method will be $\mathcal{E}(\delta,\Delta t)=\mathcal{O}(C\delta+\tilde{C}\Delta t^{p})$. Here $C$ and $\tilde{C}$ are constants that can depend on the solution and the choice of numerical method but not on $\delta$ and $\Delta t$. Here the goal is to determine the control accurately and it is therefore important that $\tilde{C}\Delta t^{p}\ll C\delta$, so that the optimization landscape depends more heavily on the real, physical change in the state vector due to the perturbation $\delta$ than on the non-physical change due to the truncation error. Clearly this is easier to achieve for a high-order method.
We note, however, that the error in the infidelity scales quadratically with the state error. Therefore, an error $\mathcal{E}(\delta,\Delta t)$ in the state vector causes an error $\mathcal{E_{\mathcal{I}}}(\delta,\Delta t)$ in the infidelity which is $\mathcal{O}(\mathcal{E}(\delta,\Delta t)^{2})$. As a result, in order to accurately optimize the infidelity up to a tolerance of $\eta$, $\Delta t$ should be chosen so that the truncation error $\tilde{C}\Delta t^{p}$ is on the order of $\sqrt{\eta}$ or smaller. For other objective functions, the error may scale linearly with the state error. In that case, a higher level of accuracy in the state will be required.
@@
As an example, we optimize the generalized infidelity of a CNOT gate for a system of two qubits coupled to a resonator bus (which will be described in more detail in Section 4.2). The optimization is performed using a fourth-order method to solve Schrödinger’s equation numerically, with the timestep size chosen so that the relative final state error is approximately $10^{-1}$. For each optimization iteration, we solve the same equation again, choosing the timestep size so that the relative final state error is approximately $10^{-7}$. In Figure 2, we plot the generalized infidelity as computed by the low-accuracy and high-accuracy solutions, as well as the relative final state error in the low-accuracy solution, as a function of the number of optimization iterations. We see that although after 500 iterations the optimization produces a gate with a generalized infidelity of approximately $10^{-4}$ as computed by the low-accuracy method, the actual generalized infidelity is closer to $10^{-1}$, and the optimization iterations performed after the generalized infidelity had reached $10^{-2}$ were wasted computations which actually increased the generalized infidelity. Comparing the high-accuracy infidelity to the final state relative error of the low-accuracy method, we see that the infidelity is generally bounded below by the square of the state error, which is consistent with our above prediction that $\Delta t$ should be chosen so that the truncation error is on the order of $\sqrt{\eta}$, where $\eta$ is the desired tolerance in the infidelity.
Figure 2: Generalized infidelity of a CNOT gate, as computed by a low-accuracy and high-accuracy numerical method. The generalized infidelity is optimized using a small number of timesteps. For each set of control parameters found at each iteration of the optimization, we also simulate the system using these control parameters and a larger number of timesteps in order to obtain a high-accuracy solution, which is used to compute a more accurate generalized infidelity and the error in the low-accuracy solution at the final time.